4. Problems and Applications Q6 The following table describes the production possibilities of two cities in the country of Baseballia: Pairs of Red Socks per Worker per Hour Pairs of White Socks per Worker per Hour Boston 2 Chicago 4 Without trade, the price of a pair of white socks (in terms of red socks) in Boston is 1 1/3 pairs ▼ of red socks, and in Chicago it is 2 1/2 pairsof red socks. Boston ▼ has an absolute advantage in the production of red socks, and Chicago ▼ has an absolute advantage in the production of white socks Boston ▼ has a comparative advantage in the production of red socks, and Chicago ▼ has a comparative advantage in the production of white socks If the cities trade with each other, Boston will export red ▼ socks, and Chicago will export white ▼ socks The price of white socks can be expressed in terms of red socks. The highest price at which white socks can be traded that would make both cities better off is 3/4 pair ▼ of red socks per pair of white socks, and the lowest price that makes both cities better off is 2/5 pair of red socks per pair of white socks. Grade It Now Save & Continue

can you please check my answers and correct if nessacery
Answer #1

Price of a pair of white socks in Boston = 5/2 = 2 1/2 pairs of red socks

Price of a pair of white socks in Chicago= 4/3 = 1 1/3 pairs of red socks

Blanks concerning absolute and comparative advantage are correct

Boston will export red socks and Chicago will export white socks.

Lowest price = 4/3 pairs of red socks and Highest price = 5/2 pairs of red socks.